# Quantum nonlinear effects in the number-conserving analogue gravity of Bose-Einstein condensates

### v1

@@ III.5 Dispersion relation
To understand the nature of the dispersion relation, let us consider a plane wave solution for (35)
@@
where anticipating possible dissipative effects, the frequency $\omega$ can in general be complex. Substituting this ansatz back into Eq. (32), and assuming that spatiotemporal variations of the background fields are negligible, we have
$\omega=v_{b}k+\frac{1}{2}\Big{(}i(\ddot{f}_{1}-f^{\prime\prime}_{3})\\ \pm\sqrt{-(\ddot{f}_{1}-f^{\prime\prime}_{3})^{2}+4(\tilde{U}f^{\prime\prime}_{1}-\ddot{f}_{1}f^{\prime\prime}_{3}+c^{2}_{s}k^{2})}\Big{)}.$ (45)
Here, we have defined the effective sound velocity of the background ($\rho_{b}$, $\theta_{b}$) to be $c_{s}^{2}=\frac{\tilde{U}\rho_{b}}{m}$. As expected, when $f_{1}=f_{3}=0$, this dispersion relation reduces to the standard Bogoliubov dispersion in the hydrodynamic limit. The appearance of dissipation, signified by a complex frequency, is a direct manifestation of the backreaction of the depleted cloud, causing the condensate particle number to be not separately conserved. The exact rate at which this occurs, and the corresponding correction on the leading-order Bogoliubov dispersion depends on the relevant field configurations leading to the solutions of the BdG equation (6).
In the Appendix, we discuss a specific time dependent configuration for a BEC contained in a box potential, where the quantization of the depletion field $\hat{\chi}$ can be performed analytically [35]. Consequently, it is possible to calculate the anomalous correlators appearing in the definitions of $f_{1}$ and $f_{3}$. As we explain in the Appendix in detail, it can be concluded that for early times, after switching on interactions from an initially noninteracting state, the particle loss from the condensate, quantified by the function $f_{1}$, is quadratic in time, and this loss is in such a setup generic in the sense that it is independent of the system size, at least for sufficiently early times. As a result, the leading order correction to the Unruh paradigm metric is also generic and governed by $h_{\mu\nu}$ in (43), until the Bogoliubov approximation breaks down (up to the order considered).
@@ IV.1 Metric coupled to second-order fluctuations
$\mathbf{g}_{tt}=g_{tt}-\sqrt{\frac{\rho_{0}}{mU}}\Big{[}\big{(}\epsilon U\frac{\delta\rho_{1}}{m}-2\epsilon v_{0}v_{1}-\epsilon^{2}v^{2}_{1}\big{)}\\ +\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\big{(}U\frac{\rho_{0}}{m}-v^{2}_{0}+\epsilon U\frac{\delta\rho_{1}}{m}-2\epsilon v_{0}v_{1}\big{)}-\epsilon^{2}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}\big{(}U\frac{\rho_{0}}{m}-v^{2}_{0}\big{)}\Big{]},$ (46)
$\mathbf{g}_{ti}=g_{ti}-\sqrt{\frac{\rho_{0}}{mU}}\Big{[}\epsilon v_{1(i)}+\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\big{(}v_{0}+\epsilon v_{1}\big{)}_{i}-\epsilon^{2}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}v_{0(i)}\Big{]}.$ (47)
$\mathbf{g}_{ij}=g_{ij}+\sqrt{\frac{\rho_{0}}{mU}}\Big{[}\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}.$ (48)
Here, we have defined the first-order correction to the velocity field as $v_{1}=\frac{\nabla\delta\theta_{1}}{m}$, and retained terms up to $\mathcal{O}(\epsilon^{2})$ in the first-order perturbed fields $\delta\rho_{1}$ and $\delta\theta_{1}$, which represents the working order for the new background-II.
The standard linearized Unruh paradigm has the metric components
$g_{tt}=-\sqrt{\frac{\rho_{0}}{mU}}\left(\frac{U\rho_{0}}{m}-v_{0}^{2}\right),\qquad g_{ti}=-\sqrt{\frac{\rho_{0}}{mU}}(v_{0})_{i},\\ g_{ij}=\sqrt{\frac{\rho_{0}}{mU}}\delta_{ij}.$ (49)
@@
where $h_{\mu\nu}$ can be thought of as emergent, as it is due to the inherent nonlinearity of background-II. From Eq. (43), we conclude that the leading-order corrections come from the nonlinear contribution to the modified metric ($h_{\mu\nu}$) with respect to the background-II. In fact, as might be expected on general physical grounds, this correction is of the same order as the density of the depleted particles $\rho_{\hat{\chi}}=[\mathcal{O}(N^{0})]$.
@@ IV.2 Dynamics of the background field generating curved spacetime
To understand how changing the form of background affects the analogue gravity description in our formalism, we start with a simple example. Let us consider a uniform and static fluid medium in one spatial dimension, described by the variables $\rho_{0},\theta_{0}$. Then the metric for linear sound propagation is Minkowski, which corresponds to background-I. Now we consider the propagation of a nonlinear perturbation; then, as we described in the above, it is possible to obtain a Lorentz-invariant description only when we couple the second-order field $\delta\theta_{2}$ with the background-II, which now however does not represent a simple Minkowski spacetime. The components of the tensor $h_{\mu\nu}$ in this case can be obtained by setting $v_{0}=0$ and $\rho_{0}=$ constant in Eqs. (46), (48), and (47). It is also important to note that we have imposed an extra condition on the fluid variables, namely that the initial configuration $\rho_{0},\theta_{0}$ satisfies the fluid-dynamical equations separately, which however is not necessary to obtain the metric for second-order fluctuation fields. This relates to employing distinct boundary and initial conditions, which are imposed on the quasi-linear fluid dynamical equations we are dealing with, and again emphasizes how the distinction between the background and the fluctuation fields is to some extent arbitrary and can thus also change an analogue gravity description significantly.
@@
$\begin{split}\partial_{t}\Big{[}-\frac{1}{\lambda}\big{(}\partial_{t}\delta\theta_{1}+\frac{1}{2}v_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2m}(\nabla\delta\theta_{1})^{2}+\frac{\partial f_{3}}{\partial\theta_{c}}|_{\theta_{0}}\delta\theta_{1}\big{)}\Big{]}+\nabla\cdot\Big{[}\rho_{0}\frac{\nabla\delta\theta_{1}}{m}-\frac{v_{0}}{\lambda}\Big{(}\partial_{t}\delta\theta_{1}+\frac{1}{2}v_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2m}(\nabla\delta\theta_{1})^{2}\\ +\frac{\partial f_{3}}{\partial\theta_{c}}|_{\theta_{0}}\delta\theta_{1}\Big{)}-\frac{\epsilon}{m\lambda}(\nabla\delta\theta_{1})\Big{(}\partial_{t}\delta\theta_{1}+\frac{1}{2}v_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2m}(\nabla\delta\theta_{1})^{2}+\frac{\partial f_{3}}{\partial\theta_{c}}|_{\theta_{0}}\delta\theta_{1}\Big{)}\Big{]}+\\ \frac{1}{\lambda}\frac{\partial f_{1}}{\partial\rho_{c}}|_{\rho_{0}}\Big{(}\partial_{t}\delta\theta_{1}+\frac{1}{2}v_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2m}(\nabla\delta\theta_{1})^{2}+\frac{\partial f_{3}}{\partial\theta_{c}}|_{\theta_{0}}\delta\theta_{1}\Big{)}-\frac{\partial f_{1}}{\partial\theta_{c}}|_{\theta_{0}}\delta\theta_{1}=0.\end{split}\vspace{-6em}$ (51)
We have defined the first-order modification to the coupling constant $U$,
$\lambda=U+\frac{\partial f_{3}}{\partial\rho_{c}}|_{\rho_{0}},$ (52)
which should be compared with its second-order counterpart in Eq. (34).
@@ V.1 Friedmann-Lemaître-Robertson-Walker background
In this section we will discuss how the quantum nonlinear effects coming from the backreaction of the quantized field $\hat{\chi}$ change the leading-order Unruh paradigm metric, which governs perturbations to linear order. To exemplify our arguments in the preceding sections we will consider a solution of the GPE for $\phi_{0}$, such that the leading-order $\mathcal{O}(N)$ density $\rho_{0}$ is constant and the condensate is at rest. This produces a simple analogue of the Friedmann-Lemaître-Robertson-Walker (FLRW) metric of cosmology when $c_{s0}=c_{s0}(t)$, cf. Ref. [48], where the time dependence of the sound speed is implemented by using a Feshbach resonance sweep [43] such that $U=U(t)$, with the constant density enforced by a hard-walled trap of infinite height.
Then, in accordance with the notation used in this paper, the form of the Unruh paradigm metric, as seen by the first order fluctuation field $\delta\rho_{1},\delta\theta_{1}$, is given by
$g^{\rm FLRW}_{\mu\nu}=\sqrt{\frac{\rho_{0}}{mU(t)}}\begin{bmatrix}-\frac{U(t)\rho_{0}}{m}&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix},$ (53)
@@ V.2 Renormalized FLRW Unruh metric
$\mathbf{g}_{tt}=g^{\rm FLRW}_{tt}-\sqrt{\frac{\rho_{0}}{mU(t)}}\Big{[}\Big{(}\epsilon U(t)\frac{\delta\rho_{1}}{m}-\epsilon^{2}v^{2}_{1}\Big{)}\\ +\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\Big{(}U(t)\frac{\rho_{0}}{m}+\epsilon U(t)\frac{\delta\rho_{1}}{m}\Big{)}-\epsilon^{2}U(t)\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}\frac{\rho_{0}}{m}\Big{]}\\ =-\sqrt{\frac{\rho_{0}}{mU(t)}}\Big{[}U(t)\frac{\rho_{0}}{m}+\frac{3}{2}\epsilon U(t)\frac{\delta\rho_{1}}{m}+\frac{3}{8}\epsilon^{2}U(t)\frac{\delta\rho^{2}_{1}}{8m\rho_{0}}\Big{]},$ (54)
$\mathbf{g}_{ti}=g^{\rm FLRW}_{ti}-\sqrt{\frac{\rho_{0}}{mU(t)}}\Big{[}\epsilon v_{1(i)}+\epsilon^{2}\frac{\delta\rho_{1}}{2\rho_{0}}v_{1(i)}\Big{]},$ (55)
$\begin{split}\mathbf{g}_{ij}=g^{\rm FLRW}_{ij}+\sqrt{\frac{\rho_{0}}{mU(t)}}\Big{[}\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}\\ =\sqrt{\frac{\rho_{0}}{mU(t)}}\Big{[}1+\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}.\end{split}$ (56)
Here the boundary conditions are chosen such that the background-I ($\rho_{0},v_{0}$) is a solution of the Euler and continuity equations for these variables. The dynamics of the background field $\delta\theta_{1}$ is governed by the nonlinear equation (51), which here takes the following form
$\begin{split}\partial_{t}\Big{[}-\frac{1}{\lambda(t)}\big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2m}(\nabla\delta\theta_{1})^{2}+\frac{\partial f_{3}}{\partial\theta_{c}}|_{\theta_{0}}\delta\theta_{1}\big{)}\Big{]}+\nabla\cdot\Big{[}\rho_{0}\frac{\nabla\delta\theta_{1}}{m}-\frac{\epsilon}{m\lambda(t)}(\nabla\delta\theta_{1})\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2m}(\nabla\delta\theta_{1})^{2}+\frac{\partial f_{3}}{\partial\theta_{c}}|_{\theta_{0}}\delta\theta_{1}\Big{)}\Big{]}\\ +\frac{1}{\lambda(t)}\frac{\partial f_{1}}{\partial\rho_{c}}|_{\rho_{0}}\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2m}(\nabla\delta\theta_{1})^{2}+\frac{\partial f_{3}}{\partial\theta_{c}}|_{\theta_{0}}\delta\theta_{1}\Big{)}-\frac{\partial f_{1}}{\partial\theta_{c}}|_{\theta_{0}}\delta\theta_{1}=0.\end{split}$ (57)
Solving this nonlinear partial differential equation for a given $U(t)$, even perturbatively and for the simplest case of instantaneously switching on $U$ where it simplifies considerably, and proving the convergence of the solution is a nontrivial task. We defer this to a future study which will then employ, e.g., the Poincaré-Lighthill-Kuo method of solving certain classes of partial differential equations [47]. We can however conclude already from the level of complexity of Eq. (57) that the renormalized metric will show intricate features not present in the bare FLRW metric.

### v2

@@ III.5 Dispersion relation
To understand the nature of the dispersion relation, let us consider a plane wave propagating in $x$ direction (37)
@@
where the frequency $\omega$ can in general be complex. Substituting this ansatz back into Eq. (34), and assuming that spatiotemporal variations of the background are negligible, we have
$\omega=v_{bx}k+\frac{1}{2}\left[i(\ddot{f}_{1}-f^{\prime\prime}_{3})\right.\\ \left.\pm\sqrt{4(\tilde{U}f^{\prime\prime}_{1}-\ddot{f}_{1}f^{\prime\prime}_{3}+c^{2}_{s,b}k^{2})-(\ddot{f}_{1}-f^{\prime\prime}_{3})^{2}}\right].$ (47)
Here, we have defined the effective sound velocity of the background ($\rho_{b}$, $\theta_{b}$) to be $c_{s,b}^{2}=\tilde{U}\rho_{b}$. As expected, when $f_{1}=f_{3}=0$, this dispersion relation reduces to the standard Bogoliubov dispersion. The appearance of a generally complex frequency is a direct manifestation of backreaction, causing the condensate particle number to be not separately conserved. The exact rate at which this occurs, and the corresponding correction on the leading-order Bogoliubov dispersion depends on the relevant field configurations leading to the solutions of the BdG equation (6).
In the Appendix, we discuss a specific time dependent configuration for a BEC contained in a box potential, where the quantization of the depletion field $\hat{\chi}$ can be performed analytically [36]. Consequently, it is possible to calculate the anomalous correlators appearing in the definitions of $f_{1}$ and $f_{3}$. As we explain in the Appendix in detail, it can be concluded that for early times, after switching on interactions from an initially noninteracting state, the particle loss from the condensate, quantified by the function $f_{1}$, is quadratic in time, and this loss is in such a setup generic in the sense that it is independent of the system size, at least for sufficiently early times. As a result, the leading order correction to the Unruh paradigm metric is also generic and governed by $h_{\mu\nu}$ in (45), until the Bogoliubov approximation itself breaks down.
@@ IV.1 Metric coupled to second-order fluctuations
$\mathbf{g}_{tt}=g_{tt}-\sqrt{\frac{\rho_{0}}{U}}\Big{[}\epsilon U\delta\rho_{1}-2\epsilon{\bm{v}}_{0}\cdot{\bm{v}}_{1}-\epsilon^{2}{\bm{v}_{1}}^{2}\\ +\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\big{(}c_{s0}^{2}-{\bm{v}}^{2}_{0}+\epsilon U\delta\rho_{1}-2\epsilon{\bm{v}}_{0}\cdot{\bm{v}}_{1}\big{)}-\epsilon^{2}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}\big{(}c_{s0}^{2}-{\bm{v}}^{2}_{0}\big{)}\Big{]},$ (48)
$\mathbf{g}_{ti}=g_{ti}-\sqrt{\frac{\rho_{0}}{U}}\Big{[}\epsilon{\bm{v}}_{1i}+\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\big{(}{\bm{v}}_{0}+\epsilon{\bm{v}}_{1}\big{)}_{i}-\epsilon^{2}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}{\bm{v}}_{0i}\Big{]}.$ (49)
$\mathbf{g}_{ij}=g_{ij}+\sqrt{\frac{\rho_{0}}{U}}\Big{[}\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}.$ (50)
Here, we have defined the first-order correction to the velocity field as ${\bm{v}}_{1}=\nabla\delta\theta_{1}$, and retained terms up to $\mathcal{O}(\epsilon^{2})$ in the first-order perturbed fields $\delta\rho_{1}$ and $\delta\theta_{1}$, which represents the working order for the new background-II.
The standard linearized Unruh paradigm metric has components
$g_{tt}=-\sqrt{\frac{\rho_{0}}{U}}\left(c_{s0}^{2}-{\bm{v}}_{0}^{2}\right),\qquad g_{ti}=-\sqrt{\frac{\rho_{0}}{U}}{\bm{v}}_{0i},\\ g_{ij}=\sqrt{\frac{\rho_{0}}{U}}\delta_{ij}.$ (51)
@@
where $h_{\mu\nu}$ can be thought of as emergent, as it is due to the inherent nonlinearity of background-II. From Eq. (45), we conclude that the leading-order corrections come from the nonlinear contribution to the modified metric ($h_{\mu\nu}$) with respect to the background-II. In fact, as might be expected on general physical grounds, this correction is of the same order as the density of the depletion cloud $\rho_{\hat{\chi}}=[\mathcal{O}(N^{0})]$.
@@ IV.2 Dynamics of the background field generating curved spacetime
To understand how changing the form of background affects the analogue gravity description in our formalism, we start with a simple example. Let us consider a uniform and static fluid medium in one spatial dimension, described by the variables $\rho_{0},\theta_{0}$. Then the metric for linear sound propagation is Minkowski, which corresponds to background-I. Now we consider the propagation of a nonlinear perturbation; then, as we described in the above, it is possible to obtain a Lorentz-invariant description only when we couple the second-order field $\delta\theta_{2}$ with the background-II, which now however does not represent a simple Minkowski spacetime. The components of the tensor $h_{\mu\nu}$ in this case can be obtained by setting ${\bm{v}}_{0}=0$ and $\rho_{0}=$ constant in Eqs. (48), (50), and (49). It is also important to note that we have imposed an extra condition on the fluid variables, namely that the initial configuration $\rho_{0},\theta_{0}$ satisfies the fluid-dynamical equations separately, which however is not necessary to obtain the metric for second-order fluctuation fields. This relates to employing distinct boundary and initial conditions, which are imposed on the quasi-linear fluid dynamical equations we are dealing with, and again emphasizes how the distinction between the background and the fluctuation fields is to some extent arbitrary and can thus also change an analogue gravity description significantly.
@@
$\begin{split}\partial_{t}\Big{[}-\frac{1}{\lambda}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\Big{]}+\nabla\cdot\Big{[}\rho_{0}\nabla\delta\theta_{1}-\frac{{\bm{v}}_{0}}{\lambda}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\\ -\frac{\epsilon}{\lambda}(\nabla\delta\theta_{1})\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\Big{]}\\ +\frac{1}{\lambda}\left.\frac{\partial f_{1}}{\partial\rho_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}-\left.\frac{\partial f_{1}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\\ -\frac{\epsilon}{\lambda}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+{\bm{v}}_{0}\cdot\nabla\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\delta\theta_{1}=0.\end{split}\vspace{-6em}$ (53)
We have defined the first-order modification to $U$ as
$\lambda=U+\left.\frac{\partial f_{3}}{\partial\rho_{c}}\right|_{0}$ (54)
neglecting a sub-leading correction of $\mathcal{O}(\epsilon^{2})$, which should be compared with its second-order counterpart in Eq. (36).
@@ V.1 Friedmann-Lemaître-Robertson-Walker background
We will now discuss in more detail how the quantum nonlinear effects coming from the backreaction of the quantized field $\hat{\chi}$ change the leading-order Unruh paradigm metric, which governs perturbations to linear order. To exemplify our arguments in the preceding sections we will consider a solution of the GPE for $\phi_{0}$, such that the leading-order $\mathcal{O}(N)$ density $\rho_{0}$ is constant and the condensate is at rest. This produces a simple analogue of the Friedmann-Lemaître-Robertson-Walker (FLRW) metric of cosmology when $c_{s0}=c_{s0}(t)$, cf. Ref. [51], where the time dependence of the sound speed is implemented by using a Feshbach resonance sweep [44] such that $U=U(t)$, with the constant density enforced by a hard-walled trap of infinite height.
Then the Unruh paradigm metric, as seen by the first order fluctuation field $\delta\rho_{1},\delta\theta_{1}$, is given by
$g^{\rm FLRW}_{\mu\nu}=\sqrt{\frac{\rho_{0}}{U(t)}}\begin{bmatrix}-{c_{s0}^{2}(t)}&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix},$ (55)
@@ V.2 Renormalized FLRW Unruh metric
$\mathbf{g}_{tt}=g^{\rm FLRW}_{tt}-\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}\Big{(}\epsilon U(t)\delta\rho_{1}-\epsilon^{2}{\bm{v}}^{2}_{1}\Big{)}\\ +\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\Big{(}U(t)\rho_{0}+\epsilon U(t)\delta\rho_{1}\Big{)}-\epsilon^{2}U(t)\rho_{0}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}\Big{]}\\ =-\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}U(t)\rho_{0}+\frac{3}{2}\epsilon U(t)\delta\rho_{1}+\frac{3}{8}\epsilon^{2}U(t)\frac{\delta\rho^{2}_{1}}{8\rho_{0}}\Big{]},$ (56)
$\mathbf{g}_{ti}$ $=$ $g^{\rm FLRW}_{ti}-\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}\epsilon{\bm{v}}_{1i}+\epsilon^{2}\frac{\delta\rho_{1}}{2\rho_{0}}{\bm{v}}_{1i}\Big{]},$ (57)
$\mathbf{g}_{ij}$ $=$ $g^{\rm FLRW}_{ij}+\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}$ (58)
$=$ $\sqrt{\frac{\rho_{0}}{U(t)}}\Big{[}1+\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}.$
Here the boundary conditions are chosen such that the background-I ($\rho_{0},{\bm{v}}_{0}$) is a solution of the Euler and continuity equations for these variables. The dynamics of the background field $\delta\theta_{1}$ is governed by the nonlinear equation (53), which here takes the somewhat simplified but still rather involved form
$\begin{split}\partial_{t}\Big{[}-\frac{1}{\lambda(t)}\big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\big{)}\Big{]}+\nabla\cdot\Big{[}\rho_{0}\nabla\delta\theta_{1}-\frac{\epsilon}{\lambda(t)}(\nabla\delta\theta_{1})\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\Big{]}\\ +\frac{1}{\lambda(t)}\left.\frac{\partial f_{1}}{\partial\rho_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}-\left.\frac{\partial f_{1}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}-\frac{\epsilon}{\lambda(t)}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\Big{(}\partial_{t}\delta\theta_{1}+\frac{\epsilon}{2}(\nabla\delta\theta_{1})^{2}+\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}\Big{)}\delta\theta_{1}=0.\end{split}$ (59)
Solving this nonlinear partial differential equation for a given $U(t)$, even for the simplest case of instantaneously switching from zero to a finite $U$, where it simplifies considerably, and proving the convergence of the solution is a nontrivial task. We defer this to a future study. Solving similar classes of partial differential equations is performed for example by the Poincaré-Lighthill-Kuo method [50]. We can however conclude already from the level of complexity of Eq. (59) that the renormalized metric will show features not present in the bare FLRW metric.